OpenStudy (anonymous):

Please i need someone to help me with a couple questions ! medal ! There are 5 positions available in the new school. Of the applicants, 12 are men and 8 are women. In how many ways can 3 men and 2 women be chosen if they are equally considered. Select one: a. 6,160 b. 220 c. 28 d. 3,375

OpenStudy (anonymous):

Evaluate: 10 C 2 Select one: a. 45 b. 90 c. 120 d. 180

OpenStudy (anonymous):

From a band of 16 members, find the total number of different arrangements of 9 musicians. Select one: a. C(16, 9) = 160 b. C(16, 9) = 11,440 c. P(16, 9) = 5,040 d. P(16, 9) = 362,880

OpenStudy (anonymous):

Which of the following is a permutation? Select one: a. A selection of 3 candy bars from 5 chocolate bars. b. First place and runner up for Miss America. c. A hand of 8 cards from a deck. d. Selecting 10 students from the tenth grade for a field trip.

OpenStudy (anonymous):

Evaluate: 6 P 4 Select one: a. 72 b. 24 c. 360 d. 2

OpenStudy (anonymous):

In how many ways can a committee of 5 be chosen from 10 people? Select one: a. 252 b. 2,002 c. 30,240 d. 100,000

OpenStudy (anonymous):

There are 5 positions available in the new school. Of the applicants, 12 are men and 8 are women. What is the probability that 5 men are selected if they are randomly picked? Select one: a. 33/646 b.5/12 c.4/5 d.9/70

OpenStudy (kropot72):

\[\large P(5\ men)=\frac{12C5}{20C5}=you\ can\ calculate\]

OpenStudy (lynfran):

Anwers (except the first one) 10C2=45 (2nd one) 11440 (3rd one) nPr=360 (4th one) nCr=252 (5th one) 33/646 (last one)

OpenStudy (anonymous):

The upward velocity, v, of the water in a fountain is 28 more than the product of -32 and t, where t is the number of seconds after the water leaves the fountain. As the water moves upward, its velocity decreases until it reaches the top where the velocity is zero. A. Write an equation to represent this. B. How long does it take for the water to reach maximum height? (v = 0)

OpenStudy (anonymous):

can you help with this please please !?

OpenStudy (anonymous):

1. The perimeter of a rectangle is 78 cm. The length is 3 greater than twice its width. Find the dimensions.

OpenStudy (anonymous):

One angle of a triangle is 2 more than four times as large as the other angle. The third angle is three times the first angle decreased by 6. What is the measure of each angle? first angle second angle third angle

OpenStudy (anonymous):

for the perimeter one divide by 2 to get 39 then since a number is 3 times greater than a different number doubled do 3*2=6 so its six times greater

OpenStudy (anonymous):

thanks !